# Field-orientation-dependent magnetic phases in GdRu2Si2 probed with muon-spin spectroscopy

### v2

@@ III.1 TF $\bm{\mu}^{+}$ measurements
We now turn to the spectra measured at $T=5$ K, where several distinct magnetic structures exist as a function of applied magnetic field [5, 7]. FT $\mu^{+}$ spectra are shown in Fig. 3 for each orientation. For $B<2.25$ T, the spectra comprise a split peak, close to the applied field. For $\bm{B}_{0}\parallel$ [001] [Figs. 3(b-d)], the spectra contain additional spectral weight at fields below the applied field, whereas for $\bm{B}_{0}\parallel$ [100], we see a pair of satellite peaks, above and below $B_{0}$. As before, this reflects the symmetry of the muon site for each field orientation. However, the features in the spectra are much broader for $T=5$ K, suggesting that the system adopts magnetic structures that give rise to a wider distribution of fields. These become sharper at higher applied fields, as the field-polarized component of the magnetic structure becomes more significant.
@@
The 2.25 T spectrum is qualitatively different from those measured at 2.5 [Fig. 3(c)] and 4 T [Fig. 3(d)] for this orientation, allowing us to clearly distinguish the response of the muon to the SkL from that arising from the other multi-Q magnetic structures that are realized in this system. For the higher fields along [001], this has been proposed to be the checkerboard antiferroic order of meron/anti-meron-like textures (MLc) [7]. The $B_{0}\geq 2.5$ T spectra both comprise a satellite peak below the applied field, although the shape of the satellite is distinct in each case. Our simulated spectra based on the MLc magnetic structure show good agreement with experiment for both values of $B_{0}$. The difference between the spectra measured at 2.25 [Fig. 3(b)] and 2.5 T [Fig. 3(c)] can be explained by the fact that the helical components of the SkL magnetic structure rotate into and out of the $c$-axis direction, either aligning or antialigning with the ferromagnetic component, whereas the modulated parts of the moment in the MLc phase remain within the $ab$ plane (and therefore perpendicular to the ferromagnetic component). Thus the former magnetic structure more effectively splits the spectral weight into distinct peaks, whereas the latter only results in broadening of the low-field feature.
@@
Turning to the spectra measured for $\bm{B}_{0}\parallel$ [100], we note that, in contrast to $\bm{B}_{0}\parallel$ [001], applied fields with magnitudes $B_{0}=2.25$ T and $B_{0}=2.5$ T in this orientation are expected to result in a single-Q conical magnetic structure. The spectra measured in this phase [Fig. 3(f) and Fig. 3(g)] comprise pairs of satellite peaks, one above $B_{0}$ and one below. These features are fairly broad and somewhat difficult to resolve in experiment for $B_{0}=2.25$ T, but become sharper at $B_{0}=2.5$ T. Our simulated spectra can be interpreted as a pair of shifted Overhauser distributions [23] resulting from a helical magnetic structure propagating along the $a$ axis, which leads to different magnetic fields for muons sitting along the $a$ axis to those lying on the $b$ axis. An additional ferromagnetic component shifts the centers of these distributions, such that one is centered on a field above $B_{0}$ and one on a field below, and also results in a small splitting of the peaks within this distribution, which is particularly noticeable for the distribution at fields below $B_{0}$. It is difficult to reconcile these spectra with those observed in experiment. The peaks observed in the experimental spectra suggest the presence of a significant field-polarized ferromagnetic component, but these are broadened, rather than experiencing the splitting that our simulations predict. We therefore suggest that these spectra are incompatible with the proposed single-Q magnetic structure. It is possible that the magnetic structure for this phase needs to be revised, and further work on this phase would be instructive.
@@
III.2 ZF $\bm{\mu}^{+}$ measurements
Zero-field $\mu^{+}$ measurements were made on two aligned GdRu2Si2 single crystals. The ZF $\mu^{+}$ spectra exhibit a very quickly relaxing oscillating asymmetry. The first 0.1 $\mu$s of the spectra were fitted to the function
@@ III.2 ZF $\bm{\mu}^{+}$ measurements
The extracted fluctuation rates $\nu\approx$3 MHz are significantly smaller than the relaxation rates $\lambda_{1}$ that accompany the relaxing part of the asymmetry, indicating that this relaxation is mostly due to the distribution of static magnetic fields resulting from the incommensurate magnetic structure. The discontinuous behavior of $\lambda_{1}$ at $T_{\mathrm{N1}}$ is therefore evidence for a significant change in the magnetic structure at this temperature. The fluctuation rates $\nu$ obtained from our simulations are similar in magnitude to the slow relaxation rate $\lambda_{2}$, consistent with the interpretation that this relaxation is due to dynamics flipping the muon spin at sites where the local magnetic field is parallel to the muon spin, in the slow fluctuation limit. We note that the relative magnitudes of $A_{1}$ and $A_{2}$ are consistent with those of the oscillating and non-oscillating parts, respectively, of the polarization function obtained by our model.
Unlike the magnetic ground state [8], the higher-temperature magnetically ordered phase has not been unambiguously identified. However, a recent powder neutron powder diffraction study [25] proposed three candidate single-q magnetic structures for this phase, with these measurements being unable to distinguish between these. We therefore simulated $\mu^{+}$ spectra based on each of the models and fit these to the data measured at $T\geq 38$ K. All of these models provided similar quality fits to the data, with a so-called sine magnetic structure with moments along the [011] direction being the best, and also being the only model that gives a magnetic moment that decreases monotonically with temperature, as shown in Fig. 5(d). (We note that for these sine magnetic structures, which do not have a constant magnetic moment amplitude, $\mu_{\mathrm{Gd}}$ represents the maximum value of the magnetic moment.) This magnetic structure also provides a better fit to the spectrum measured at $T=38$ K than the constant-moment solution, as indicated by the blue line in Fig. 5(c). Our modelling also indicates that the fluctuation rate increases on approaching $T_{\mathrm{N2}}$ from below, as shown in Fig. 5(e), but remains within the slow fluctuation limit. This suggests that the distribution of static magnetic fields resulting from an incommensurate magnetic structure is also responsible for the majority of the relaxation observed in this higher-temperature ordered phase.
@@ IV Conclusion
We have illustrated how implanted muons are sensitive to the magnetic phases in GdRu2Si2, providing evidence for skyrmion and meron-lattice magnetic phases. In contrast, $\mu^{+}$ spectra measured for fields along [100] are inconsistent with the previously proposed single-Q magnetic structure, suggesting that this phase needs to be revised. We demonstrate a significant, negative hyperfine contribution to the local magnetic field at the muon site, due to conduction electrons becoming spin-polarized through the RKKY interaction. Zero field measurements as a function of temperature demonstrate a transition between two distinct bulk magnetically ordered states at $T_{\mathrm{N}1}=38.8(1)$ K.
@@ Acknowledgements.
Part of the work was carried out at the Swiss Muon Source, Paul Scherrer Institute, Switzerland and we are grateful for the provision of beamtime and experimental support. We acknowledge computing resources provided by Durham Hamilton HPC. This work was funded by the EPSRC UK Skyrmion Project under Grant No. EP/N032128/1. The work at the University of Warwick was further supported by EPSRC grant EP/T005963/1. Work at Oxford was funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (Grant No. EP/X025861/1). AHM is grateful to STFC and EPSRC for the provision of a studentship under Grant No. EP/T518001. MG acknowledges the financial support of the Slovenian Research and Innovation Agency through Program No. P1-0125 and Projects No. J1-50008, J1-50012, and and N1-0356.

### v3

@@ III.2 Transverse-field $\bm{\mu}^{+}$ measurements
We now turn to the spectra measured at $T=5$ K, where several distinct magnetic structures exist as a function of applied magnetic field [5, 7]. FT $\mu^{+}$ spectra are shown in Fig. 4 for each orientation. For $B<2.25$ T, the spectra comprise a split peak, close to the applied field. For $\bm{B}_{0}\parallel$ [001] [Figs. 4(b–d)], the spectra contain additional spectral weight at fields below the applied field, whereas for $\bm{B}_{0}\parallel$ [100], we see a pair of satellite peaks, above and below $B_{0}$. As before, this reflects the symmetry of the muon site for each field orientation. However, the features in the spectra are much broader for $T=5$ K, suggesting that the system adopts magnetic structures that give rise to a wider distribution of fields. These become sharper at higher applied fields, as the field-polarized component of the magnetic structure becomes more significant.
@@
The 2.25 T spectrum is qualitatively different from those measured at 2.5 [Fig. 4(c)] and 4 T [Fig. 4(d)] for this orientation, allowing us to clearly distinguish the response of the muon to the SkL from that arising from the other multi-Q magnetic structures that are realized in this system. For the higher fields along [001], this has been proposed to be the checkerboard antiferroic order of meron/anti-meron-like textures (MLc) [7]. The $B_{0}\geq 2.5$ T spectra both comprise a satellite peak below the applied field, although the shape of the satellite is distinct in each case. Our simulated spectra based on the MLc magnetic structure show fair agreement with experiment for both values of $B_{0}$. The difference between the spectra measured at 2.25 [Fig. 4(b)] and 2.5 T [Fig. 4(c)] can be explained by the fact that the helical components of the SkL magnetic structure rotate into and out of the $c$-axis direction, either aligning or antialigning with the ferromagnetic component, whereas the modulated parts of the moment in the MLc phase remain within the $ab$ plane (and therefore perpendicular to the ferromagnetic component). Thus the former magnetic structure more effectively splits the spectral weight into distinct peaks, whereas the latter only results in broadening of the low-field feature.
@@
Turning to the spectra measured for $\bm{B}_{0}\parallel$ [100], we note that, in contrast to $\bm{B}_{0}\parallel$ [001], applied fields with magnitudes $B_{0}=2.25$ T and $B_{0}=2.5$ T in this orientation are expected to result in a single-Q conical magnetic structure. The spectra measured in this phase [Fig. 4(f) and Fig. 4(g)] comprise pairs of satellite peaks, one above $B_{0}$ and one below. These features are fairly broad and somewhat difficult to resolve in experiment for $B_{0}=2.25$ T, but become sharper at $B_{0}=2.5$ T. Our simulated spectra can be interpreted as a pair of shifted Overhauser distributions [27] resulting from a helical magnetic structure propagating along the $a$ axis, which leads to different magnetic fields for muons sitting along the $a$ axis to those lying on the $b$ axis. An additional ferromagnetic component shifts the centers of these distributions, such that one is centered on a field above $B_{0}$ and one on a field below, and also results in a small splitting of the peaks within this distribution, which is particularly noticeable for the distribution at fields below $B_{0}$. It is difficult to reconcile these spectra with those observed in experiment. The peaks observed in the experimental spectra suggest the presence of a significant field-polarized ferromagnetic component, which was included in the magnetic structure used in our simulations, but these are broadened, rather than experiencing the splitting that our simulations predict. We therefore suggest that these spectra are incompatible with the proposed single-Q magnetic structure. It is possible that the magnetic structure for this phase needs to be revised, and further work on this phase would be instructive.
@@
III.3 Zero-field $\bm{\mu}^{+}$ measurements
Zero-field $\mu^{+}$ measurements were made on two aligned GdRu2Si2 single crystals using the GPS spectrometer. The ZF $\mu^{+}$ spectra exhibit a very quickly relaxing oscillating asymmetry. The first 0.1 $\mu$s of the spectra were fitted to the function
@@ III.3 Zero-field $\bm{\mu}^{+}$ measurements
The extracted fluctuation rates $\nu\approx$ 3 MHz are significantly smaller than the relaxation rates $\lambda_{1}$ that accompany the relaxing part of the asymmetry, indicating that this relaxation is mostly due to the distribution of static magnetic fields resulting from the incommensurate magnetic structure. They are also much smaller than the precession frequency $f$, implying we are in a slowly fluctuating regime (assuming the fluctuating amplitude is of a similar order of magnitude to the static field). The discontinuous behavior of $\lambda_{1}$ at $T_{\mathrm{N1}}$ is therefore evidence for a significant change in the magnetic structure at this temperature. The fluctuation rates $\nu$ obtained from our simulations are similar in magnitude to the slow relaxation rate $\lambda_{2}$, consistent with the interpretation that this relaxation is due to dynamics flipping the muon spin at sites where the local magnetic field is parallel to the muon spin, in the slow fluctuation limit. We note that the relative magnitudes of $A_{1}$ and $A_{2}$ are consistent with those of the oscillating and non-oscillating parts, respectively, of the polarization function obtained by our model.
Unlike the magnetic ground state [8], the higher-temperature magnetically ordered phase has not been unambiguously identified. However, a recent powder neutron powder diffraction study [9] proposed three candidate single-Q magnetic structures for this phase, with these measurements being unable to distinguish between these. We therefore simulated $\mu^{+}$ spectra based on each of the models and fit these to the data measured at $T\geq 38$ K. All of these models provided similar quality fits to the data, with a so-called sine magnetic structure with moments along the [011] direction being the best, and also being the only model that gives a magnetic moment that decreases monotonically with temperature, as shown in Fig. 6(d). (We note that for these sine magnetic structures, which do not have a constant magnetic moment amplitude, $\mu_{\mathrm{Gd}}$ represents the maximum value of the magnetic moment.) This magnetic structure also provides a better fit to the spectrum measured at $T=38$ K than the constant-moment solution, as indicated by the blue line in Fig. 6(c). Our modeling also indicates that the fluctuation rate increases on approaching $T_{\mathrm{N2}}$ from below, as shown in Fig. 6(e), but remains within the slow fluctuation limit. This suggests that the distribution of static magnetic fields resulting from an incommensurate magnetic structure is also responsible for the majority of the relaxation observed in this higher-temperature ordered phase.
@@ IV Conclusion
We have illustrated how implanted muons are sensitive to the magnetic phases in GdRu2Si2, providing evidence for skyrmion and meron-lattice magnetic phases. In contrast, $\mu^{+}$ spectra measured for fields along [100] are inconsistent with the previously proposed single-Q magnetic structure, suggesting that this phase needs to be revised. We demonstrate a significant, negative hyperfine contribution to the local magnetic field at the muon site, due to conduction electrons becoming spin-polarized through the RKKY interaction. This significant hyperfine field is also observed in our angle-dependent measurements of the muon Knight shift, which also provide strong experimental evidence for the muon stopping sites that form the basis of much of our analysis. Zero field measurements as a function of temperature demonstrate a transition between two distinct bulk magnetically ordered states at $T_{\mathrm{N}1}=38.8(1)$ K.
@@ Acknowledgements.
Part of the work was carried out at the Swiss Muon Source, Paul Scherrer Institute, Switzerland and we are grateful for the provision of beamtime and experimental support. We acknowledge computing resources provided by Durham Hamilton HPC. This work was funded by the EPSRC UK Skyrmion Project under Grant No. EP/N032128/1. The work at the University of Warwick was further supported by EPSRC grant EP/T005963/1. Work at Oxford was funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (Grant No. EP/X025861/1). AHM is grateful to STFC and EPSRC for the provision of a studentship under Grant No. EP/T518001. MG acknowledges the financial support of the Slovenian Research and Innovation Agency through Program No. P1-0125 and Projects No. Z1-1852, N1-0148, J1-2461, J1-50008, J1-50012, N1-0356, and N1-0345.